# Recovering contact forms from boundary data

### v5

@@ 6.1. Invariance of the volumes $\{\kappa_{j}(\beta)\}$ under special deformations of $\beta$
Let $\dim X=\dim Y=5$. We adopt the notations and hypotheses of Theorem 6.3. Then $\kappa_{1}(\beta_{X})=0=\kappa_{3}(\beta_{X})$ and
[…]
@@
Let $X$ and $(Y,\beta_{Y})$ be as in Theorem 6.3. Consider a $t$-family $\{\Psi_{t}:X\hookrightarrow\mathsf{int}(Y)\}_{t\in[0,1]}$ of orientation-preserving contact embeddings such that, for an even $j\in[2,2n]$, the $t$-family of forms $\big{\{}\Psi_{t}^{\ast}\big{(}\beta_{Y}\wedge(d\beta_{Y})^{\frac{2n-j}{2}}\big{)}\big{\}}$ is variationally boundary $(2n+1-j)$-exact in the sense of Definition 6.5.
Then $\int_{\mathcal{Y}_{j-1}^{+}(t)}(d\beta_{Y})^{\frac{2n+2-j}{2}}=-\kappa_{j}(\Psi_{t}^{\ast}\big{(}\beta_{Y}))=-\kappa^{+}_{j-1}(\Psi_{t}^{\ast}\big{(}\beta_{Y}))$ is $t$-independent. $\diamondsuit$
7. Shadows of Legandrian submanifolds of $(X,\beta)$ on the screen $\partial_{1}^{+}X(v_{\beta})$
@@
Certain properties of a contact pair $(X,\beta_{X})$ are inherited by any contact pair $(Y,\beta_{Y})$ that incapsulates $(X,\beta_{X})$ so that $\beta_{Y}|_{X}=\beta_{X}$. For example, if $(X,\beta_{X})$, $\dim X=2n+1$, contains a Legandrian sphere $S^{n}$ or a Legandrian torus $T^{n}$, then so does $(Y,\beta_{Y})$. If $(X,\beta_{X})$ contains a Legandrian sphere $S^{n}$ that bounds an embedded ball $D^{n+1}$, then so does $(Y,\beta_{Y})$. In particular, if a $3$-fold $(X,\beta_{X})$ contains a Legandrian loop $L$ that bounds a $2$-ball $B^{2}\subset X$ (such $\beta_{X}$ is called overtwisted [E]), then any contact $3$-fold $(Y,\beta_{Y})$ that incapsulates $(X,\beta_{X})$ is also overtwisted. $\diamondsuit$
Our next goal is to describe the images $L^{\dagger}$ of sub-Legendrian compact submanifolds $L\subset X$ ($\dim(L)\leq n$) both in $\partial_{1}^{+}X(v_{\beta})$ under the $(-v_{\beta})$-flow and in the trajectory space $\mathcal{T}(v_{\beta})$ under the obvious quotient map $\Gamma:X\to\mathcal{T}(v_{\beta})$. Note that these shadows of $L$ do not change when we apply a $t$-family ($t$ being sufficiently small) of local diffeomorphisms $\phi^{t}$ (as in Definition 4.4), generated by the vector field $v_{\beta}$. Since $\beta$ and $d\beta$ are invariant under $\phi^{t}$, we conclude that $\phi^{t}(L)$ is also sub-Legendrian.
@@
Let $\mathcal{X}_{2}(v_{\beta})$ denote the union of the downward $(-v_{\beta})$-trajectories through the locus $\partial_{2}X(v_{\beta})$, and $L(v_{\beta})$ the union of the downward $(-v_{\beta})$-trajectories through the points of a sub-Legandrian submanifold $L\subset X$. Put $L^{\dagger}=_{\mathsf{def}}L(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})$ (see Fig. 3).
@@
We say that a sub-Legandrian manifold $L\subset X$ is in a weakly general position to the $v_{\beta}$-flow, if $L\setminus(L\cap\mathcal{X}_{2}(v_{\beta}))$ is dense in $L$. $\diamondsuit$
Figure 3. The $(-v_{\beta})$-shadow $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ of a closed Legandrian manifold $L\subset X$, shown as a bold loop in the interior of the bulk $X$. In the figure, $L^{\dagger}$ consists of two arcs. The surface $L(v_{\beta})$ consists of $(-v_{\beta})$-trajectories falling down from $L$. Note the locus $\delta L(v_{\beta})\subset X$ that consists of $(-v_{\beta})$-trajectories falling down from $L^{\dagger}\cap\partial_{2}^{+}X(v_{\beta})$.
@@
The locus $\mathsf{int}(L^{\dagger})=L^{\dagger}\setminus(L^{\dagger}\cap\mathcal{X}_{2}(v_{\beta}))$ is an open (possibly disconnected) Lagrangian submanifold of the open $(2n)$-manifold
$\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})).$
@@
Remark 8.1. Since $L$ is embedded in $X$ and $v_{\beta}$ is transversal to $\xi$, $\mathsf{int}(L^{\dagger})$ is immersed in the open $(2n)$-manifold $\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$. $\diamondsuit$
@@
In the vicinity of $a_{\star}$, by the Darboux theorem, we can pick coordinates $(z,x_{1},y_{1},\;\ldots,\;\hfill\break x_{n},y_{n})$ such that: they vanish at $a_{\star}$, $\beta=dz+\sum_{i=1}^{n}x_{i}\,dy_{i}$, $d\beta=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}$, and $v_{\beta}=\partial_{z}$. In these coordinates, the germ at $a_{\star}$ of the Legendrian submanilold $L$ in question is given, in its parametric form, by some smooth functions $z(\vec{t}),\,x_{1}(\vec{t}),\,y_{1}(\vec{t}),\,\ldots\,,x_{n}(\vec{t}),\,y_{n}(\vec{t}),$ where $\vec{t}=(t_{1},\ldots,t_{n})\in\mathbb{R}^{n}$. We assume that the tangent to $L$ vectors
@@
By a direct computation, the second group of equations (7.2) coincides with the group of equations in (7.3). In turn, the equations in (7.3) claim that $\{d\beta(\tilde{\tau}_{i},\tilde{\tau}_{j})=0\}_{1\leq i\neq j\leq n}$. In other words, they claim that the vectors $\{\tilde{\tau}_{i}\}_{1\leq i\leq n}$ belong to some $n$-space $\mathcal{I}so(d\beta|_{H})\subset H$, so that the $n$-manifold $L_{H}^{\dagger}$, given by the functions $\{x_{1}(\vec{t}),\,y_{1}(\vec{t}),\,\ldots\,,x_{n}(\vec{t}),\,y_{n}(\vec{t})\}$ is Lagrangian in the hypersurfuce $H$. The same argument shows that the germ of any Lagrangian submanifold $L_{H}^{\dagger}\subset H$, passing through the point $\vec{0}\in H$, is the shadow, under the $z$-projection, of the germ of unique Legandrian $L$, containing the origin of $\mathbb{R}^{2n+1}$.
@@
where $C_{v_{\beta}}$ is the causality map. Note that, for a connected component $L^{\ddagger}$ of $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$, the function $\Psi^{\beta}_{L^{\ddagger}}=\Psi^{\beta}_{L^{\dagger}}\big{|}_{L^{\ddagger}}$ is continuous and smooth.
@@
Let us revisit a given Legandrian manifold $L\subset X$ whose $(-v_{\beta})$-projection is $L^{\dagger}$. For any $x\in\mathsf{int}(L^{\dagger})$, take the first point $y(x)\in\gamma_{x}$ that belongs to $L$ and consider the function
@@
where $[x,y(x)]\subset\gamma_{x}$ is the segment of the trajectory $\gamma_{x}$, bounded by $x$ and $y(x)$.
@@
Let $\beta$ be a contact form whose Reeb vector field $v_{\beta}$ is traversing and boundary generic. Let $L\subset X$ be a Legandrian submanifold in a weakly general position to “the waterfall” $\mathcal{X}_{2}(v_{\beta})$, the set of $(-v_{\beta})$-trajectories through the locus $\partial_{2}X(v_{\beta})$. Let $L^{\dagger}$ be its $(-v_{\beta})$-guided projection on $\partial_{1}^{+}X(v_{\beta})$.
@@
Then the knowledge of the shadow $L^{\dagger}$, together with the set the base points $\{b_{\kappa}\}_{\kappa}$ and the set of values $\{\psi^{\beta}_{L}(b_{\kappa})\}_{\kappa}$, makes it possible to reconstruct the Legandrian embedding $L\subset X$.
@@
Let $\pi_{v_{\beta}}:L\to L^{\dagger}$ be the $(-v_{\beta})$-guided projection. Let $L^{\dagger}_{\times}\subset L^{\dagger}$ denotes the subset consisting of points that have multiple $\pi_{v_{\beta}}$-preimages in $L$, and let $\mathcal{L}^{\dagger}_{2}=L^{\dagger}\cap\mathcal{X}_{2}(v_{\beta})$. By the hypothesis, the set $L^{\dagger}\setminus(L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2})$ is open and dense in $L^{\dagger}$.
@@
Take a point $x\in L^{\dagger}$ that belongs to the boundary of $U^{\dagger}_{\kappa}$ and such that $x\notin L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2}$. Then by Lemma 7.2, there exists a unique lift $L^{\bullet}$ to $X$ of $L^{\dagger}$ over some open neighborhood $U_{x}\subset L^{\dagger}$ of $x$, such that the point $y=\gamma_{x}\cap L$ belongs to $L^{\bullet}$. Since, over $U_{x}\cap U^{\dagger}_{\kappa}$, both $L^{\bullet}$ and $L$ are the lifts of $L^{\dagger}$ that share the point $y$, again by Lemma 7.2, we conclude that $L^{\bullet}=L$ over $U_{x}$. As a result, the reconstruction of $L$ over $U_{x}\cup U^{\dagger}_{\kappa}$ is available. Therefore, the maximal open set in $L^{\dagger}$, over which the lift of $L^{\dagger}$, subject to the constraint “$a_{\kappa}$ belongs to the lift”, is possible is a connected component $L^{\ddagger}$ of the open set $L^{\dagger}\setminus(L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2})$. Since $L$ is in weakly general position with $\mathcal{X}_{2}(v_{\beta})$, the union of such lifts, taken over all $\kappa$’s, is a dense subset $L\setminus(L\cap\pi_{v_{\beta}}^{-1}(\mathcal{L}^{\dagger}_{2}\cup L^{\dagger}_{\times}))$ of $L$. Therefore, taking the closure of this union, produces the given $L\subset X$. ∎
Remark 8.1. Note that the very existence of Legandrian $L\subset X$ insured that the assembly instructions in the proof of Corollary 7.1 were consistent.
Even when the Reeb vector field $v_{\beta}$ is convex (so the waterfall $\mathcal{X}_{2}(v_{\beta})=\partial_{2}^{-}X(v_{\beta})$) and the $v_{\beta}$-trajectories $\gamma_{x}$ depend continuously on $x\in\partial_{1}^{+}X(v_{\beta})$, we do not know whether any closed Lagrangian submanifold $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ is the $\pi_{v_{\beta}}$-shadow of a Legandrian manifold $L\subset X$. Corollary 7.1 claims only that, if such $L$ exists, then choosing a point $a_{\star}\in L$ makes it unique. It seems that to tackle this question we need to be able to estimate from above the function $\Psi^{\beta}_{L^{\dagger}}(x)$ (or at least the Reeb elevation $\mathsf{elev}_{\mathcal{R}}(L^{\dagger})$ of $L^{\dagger}$) in terms of the form $\beta$ in the vicinity of $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$. $\diamondsuit$
@@
Let $\beta$ be a contact form on $X$, and $v_{\beta}$ its traversing boundary generic Reeb vector field. Assume that Property $\mathsf{A}$ from Definition 3.1 is valid. Let $\phi:(L,\partial L)\hookrightarrow(X,\partial_{1}X)$ be a Legandrian embedding of a compact manifold $L$.
@@
  • the Lagrangian666with respect to the symplectic form $d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}$ immersion $\mathsf{int}(L^{\dagger})\subset\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$,
  • the set of base points $\{b_{\kappa}\}_{\kappa}$ and the values $\{\psi^{\beta}_{L}(b_{\kappa})\}_{\kappa}$ as in Corollary 7.1,
allow to reconstruct the Legandrian embedding $\phi:L\hookrightarrow X$, up to a diffeomorphism of $X$ that is the identity on $\partial_{1}X$.
@@
Let $X$ be a compact connected smooth $12n+$ 1-dimensional manifold, $\beta$ a contact form, and $v_{\beta}$ its Reeb vector field, which is traversing.
Let $L\subset X$ be a sub-Legandrian $(2k+1)$-dimensional closed submanifold of $X$, where $2k+1\leq n$. We assume that $L$ is in a weakly general position with respect to the waterfall $\mathcal{X}_{2}(v_{\beta})$. We denote by $L(v_{\beta})$ the set of downward trajectories originating in $L$, and by $\delta L(v_{\beta})$ the set of downward trajectories originating in $L$ and hitting the locus $\partial_{2}^{+}X(v_{\beta})$ (see Fig. 3).
Then, for any $k>0$, we get $\int_{L(v_{\beta})}(d\beta)^{k+1}=0,\;\text{ and }\;\quad\int_{L^{\dagger}}\beta\wedge(d\beta)^{k}=0.$
[…]
@@
Note that, at a generic point $x\in L(v_{\beta})$, the tangent space $T_{x}(L(v_{\beta}))$ is spanned by $v_{\beta}(x)$ and the image of a tangent space $T_{y}L$ under $(-v_{\beta})$-flow that takes $y\in L$ to $x$. Then, $(d\beta)^{k+1}|_{T_{x}(L(v_{\beta}))}=0$, since $v_{\beta}\in\ker(d\beta)$ Therefore, $\int_{L(v_{\beta})}(d\beta)^{k+1}=0$.
@@
In the special case $k=0$, using similar analysis and Stokes’ theorem, we get
[…]

### v6

@@ 6.1. Invariance of the volumes $\{\kappa_{j}(\beta)\}$ under special deformations of $\beta$
Let $\dim X=\dim Y=5$. We adopt the notations and hypotheses of Theorem 6.3. Then $\kappa_{1}(\beta_{X})=0=\kappa_{3}(\beta_{X})$ and κ_1^+(β_X) = -∫_Y_1^+(dβ_Y)^2 = κ_2(β_X),  κ_3^+(β_X) = -∫_Y_3^+ dβ_Y =κ_4(β_X).
@@
Let $X$ and $(Y,\beta_{Y})$ be as in Theorem 6.3. Consider a $t$-family $\{\Psi_{t}:X\hookrightarrow\mathsf{int}(Y)\}_{t\in[0,1]}$ of orientation-preserving contact embeddings such that, for an even $j\in[2,2n]$, the $t$-family of forms $\big\{\Psi_{t}^{\ast}\big(\beta_{Y}\wedge(d\beta_{Y})^{\frac{2n-j}{2}}\big)\big\}$ is variationally boundary (2n+1j)(2n+1-j)-exact in the sense of Definition 6.5.
Then the quantities $\int_{\mathcal{Y}_{j-1}^{+}(t)}(d\beta_{Y})^{\frac{2n+2-j}{2}}=-\kappa_{j}(\Psi_{t}^{\ast}\big(\beta_{Y}))=-\kappa^{+}_{j-1}(\Psi_{t}^{\ast}\big(\beta_{Y}))$ are $t$-indepen-dent. $\diamondsuit$
7. Shadows of Legendrian submanifolds of $(X,\beta)$ on the screen $\partial_{1}^{+}X(v_{\beta})$
@@
Certain properties of a particular contact pair $(X,\beta_{X})$ are inherited by any contact pair $(Y,\beta_{Y})$ that incapsulates $(X,\beta_{X})$ so that $\beta_{Y}|_{X}=\beta_{X}$. For example, if $(X,\beta_{X})$, $\dim X=2n+1$, contains a Legendrian sphere $S^{n}$ or a Legendrian torus $T^{n}$, then so does any ambient $(Y,\beta_{Y})$. If $(X,\beta_{X})$ contains a Legendrian sphere $S^{n}$ that bounds an embedded ball $D^{n+1}$, then so does $(Y,\beta_{Y})$. In particular, if a $3$-fold $(X,\beta_{X})$ contains a Legendrian loop $L$ that bounds an embedded $2$-ball $B^{2}\subset X$ (such $\beta_{X}$ is called overtwisted [E]), then any contact $3$-fold $(Y,\beta_{Y})$ that incapsulates $(X,\beta_{X})$ is also overtwisted. $\diamondsuit$
Our next goal is to describe the images $L^{\dagger}$ of sub-Legendrian compact submanifolds $L\subset X$ ($\dim(L)\leq n$) both in $\partial_{1}^{+}X(v_{\beta})$ under the $(-v_{\beta})$-flow projection and in the trajectory space $\mathcal{T}(v_{\beta})$ under the obvious quotient map $\Gamma:X\to\mathcal{T}(v_{\beta})$. Note that these shadows of $L$ do not change when we apply a $t$-family ($t$ being sufficiently small) of local diffeomorphisms $\phi^{t}$ (as in Definition 4.4), generated by the vector field $v_{\beta}$. Since $\beta$ and $d\beta$ are invariant under $\phi^{t}$, we conclude that $\phi^{t}(L)$ is also sub-Legendrian.
@@
Let $\mathcal{X}_{2}(v_{\beta})$ denote the union of the downward $(-v_{\beta})$-trajectories through the locus $\partial_{2}X(v_{\beta})$, and $L(v_{\beta})$ the union of the downward $(-v_{\beta})$-trajectories through the points of a sub-Legendrian submanifold $L\subset X$. Put $L^{\dagger}=_{\mathsf{def}}L(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})$ (see Fig. 3).
@@
We say that a sub-Legendrian manifold $L\subset X$ is in a weakly general position to the $v_{\beta}$-flow, if $L\setminus(L\cap\mathcal{X}_{2}(v_{\beta}))$ is dense in $L$. $\diamondsuit$
Figure 3. The $(-v_{\beta})$-shadow $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ of a closed Legendrian manifold $L\subset X$, shown as a bold loop in the interior of the bulk $X$. In the figure, $L^{\dagger}$ consists of two arcs. The surface $L(v_{\beta})$ consists of $(-v_{\beta})$-trajectories falling down from $L$. Note the locus $\delta L(v_{\beta})\subset X$ that consists of $(-v_{\beta})$-trajectories falling down from the $1$-dimensional locus $L^{\dagger}\cap\partial_{2}^{+}X(v_{\beta})$.
@@
The locus $\mathsf{int}(L^{\dagger})=_{\mathsf{def}}L^{\dagger}\setminus(L^{\dagger}\cap\mathcal{X}_{2}(v_{\beta}))$ is an open (possibly disconnected) Lagrangian submanifold of the open $(2n)$-manifold
$\partial_{1}^{+}X(v_{\beta})\setminus\big(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})\big).$
@@
Remark 8.1. Since $L$ is embedded in $X$ and $v_{\beta}$ is transversal to $\xi$, $\mathsf{int}(L^{\dagger})$ is immersed in the open $(2n)$-manifold $\partial_{1}^{+}X(v_{\beta})\setminus\big(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta})\big)$. $\diamondsuit$
@@
In the vicinity of $a_{\star}$, by the Darboux theorem, we can pick coordinates $(z,x_{1},y_{1},\;\ldots,\;\hfill\penalty-10000x_{n},y_{n})$ such that: they vanish at $a_{\star}$, and $\beta=dz+\sum_{i=1}^{n}x_{i}\,dy_{i}$, and $d\beta=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}$. Then $v_{\beta}=\partial_{z}$. In these coordinates, the germ at $a_{\star}$ of the Legendrian submanilold $L$ in question is given, in its parametric form, by some smooth functions $z(\vec{t}),\,x_{1}(\vec{t}),\,y_{1}(\vec{t}),\,\ldots\,,\hfill\penalty-10000x_{n}(\vec{t}),\,y_{n}(\vec{t}),$ where $\vec{t}=(t_{1},\ldots,t_{n})\in\mathbb{R}^{n}$. We assume that the tangent to $L$ vectors
@@
By a direct computation, the second group of equations (7.2) coincides with the group of equations in (7.3). In turn, the equations in (7.3) claim that $\{d\beta(\tilde{\tau}_{i},\tilde{\tau}_{j})=0\}_{1\leq i\neq j\leq n}$. In other words, they claim that the vectors $\{\tilde{\tau}_{i}\}_{1\leq i\leq n}$ belong to some $n$-space $\mathcal{I}so(d\beta|_{H})\subset H$, so that the $n$-manifold $L_{H}^{\dagger}$, given by the functions $\{x_{1}(\vec{t}),\,y_{1}(\vec{t}),\,\ldots\,,x_{n}(\vec{t}),\,y_{n}(\vec{t})\}$ is Lagrangian in the hypersurfuce $H$. The same argument shows that the germ of any Lagrangian submanifold $L_{H}^{\dagger}\subset H$, passing through the point $\vec{0}\in H$, is the shadow, under the $z$-projection, of the germ of unique Legendrian $L$, containing the origin of $\mathbb{R}^{2n+1}$.
@@
where $C_{v_{\beta}}$ is the causality map (see (3.1)). Note that, for a connected component $L^{\ddagger}$ of $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$, the function $\Psi^{\beta}_{L^{\ddagger}}=\Psi^{\beta}_{L^{\dagger}}\big|_{L^{\ddagger}}$ is continuous and smooth.
@@
Let us revisit the given Legendrian manifold $L\subset X$ whose $(-v_{\beta})$-projection is $L^{\dagger}$. For any $x\in\mathsf{int}(L^{\dagger})$, take the first point $y(x)\in\gamma_{x}$ that belongs to $L$ and consider the function
@@
where $x,y(x)]\subset\gamma_{x}$ is the segment of the trajectory $\gamma_{x}$, bounded by $x$ and $y(x)$.
@@
Let $\beta$ be a contact form whose Reeb vector field $v_{\beta}$ is traversing and boundary generic. Let $L\subset X$ be a Legendrian submanifold in a weakly general position to “the waterfall” $\mathcal{X}_{2}(v_{\beta})$, the set of $(-v_{\beta})$-trajectories through the locus $\partial_{2}X(v_{\beta})$. Let $L^{\dagger}$ be its $(-v_{\beta})$-guided projection on $\partial_{1}^{+}X(v_{\beta})$.
@@
Then the knowledge of the shadow $L^{\dagger}$, together with the set the base points $\{b_{\kappa}\}_{\kappa}$ and the set of values $\{\psi^{\beta}_{L}(b_{\kappa})\}_{\kappa}$, allows for a reconstruction of the Legendrian embedding $L\subset X$.
@@
Let $\pi_{v_{\beta}}:L\to L^{\dagger}$ be the $(-v_{\beta})$-guided projection. Let $L^{\dagger}_{\times}\subset L^{\dagger}$ denotes the subset consisting of points that have multiple $\pi_{v_{\beta}}$-preimages in $L$, and let $\mathcal{L}^{\dagger}_{2}=_{\mathsf{def}}L^{\dagger}\cap\mathcal{X}_{2}(v_{\beta})$. By the hypothesis, the set $L^{\dagger}\setminus(L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2})$ is open and dense in $L^{\dagger}$.
@@
Take a point $x\in L^{\dagger}$ that belongs to the boundary of $U^{\dagger}_{\kappa}$ and such that $x\notin L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2}$. Then by Lemma 7.2, there exists a unique lift $L^{\bullet}$ to $X$ of $L^{\dagger}$ over some open neighborhood $U_{x}\subset L^{\dagger}$ of $x$, such that the point $y=\gamma_{x}\cap L$ belongs to $L^{\bullet}$. Since, over $U_{x}\cap U^{\dagger}_{\kappa}$, both $L^{\bullet}$ and $L$ are the lifts of $L^{\dagger}$ that share the point $y$, again by Lemma 7.2, we conclude that $L^{\bullet}=L$ over $U_{x}$. As a result, the reconstruction of $L$ over $U_{x}\cup U^{\dagger}_{\kappa}$ is available. Therefore, the maximal open set in $L^{\dagger}$, over which the lift of $L^{\dagger}$, subject to the constraint $a_{\kappa}$ belongs to the lift”, is possible is a connected component $L^{\ddagger}$ of the open set $L^{\dagger}\setminus(L^{\dagger}_{\times}\cup\mathcal{L}^{\dagger}_{2})$. Since $L$ is in weakly general position with $\mathcal{X}_{2}(v_{\beta})$, the union of such lifts, taken over all $\kappa$’s, is a dense subset $L\setminus(L\cap\pi_{v_{\beta}}^{-1}(\mathcal{L}^{\dagger}_{2}\cup L^{\dagger}_{\times}))$ of $L$. Therefore, taking the closure of this union, produces the given $L\subset X$. ∎
Remark 7.2.
Note that the very existence of Legendrian submanifold $L\subset X$ insured that the assembly instructions in the proof of Corollary 7.1 were consistent.
Even when the Reeb vector field $v_{\beta}$ is convex (so the waterfall $\mathcal{X}_{2}(v_{\beta})=\partial_{2}^{-}X(v_{\beta})$) and the $v_{\beta}$-trajectories $\gamma_{x}$ depend continuously on $x\in\partial_{1}^{+}X(v_{\beta})$, we do not know whether any closed Lagrangian submanifold $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$ is the $\pi_{v_{\beta}}$-shadow of a Legendrian manifold $L\subset X$. Corollary 7.1 claims only that, if such $L$ exists, then choosing a point $a_{\star}\in L$ makes it unique. It seems that to tackle this question we need to be able to estimate from above the function $\Psi^{\beta}_{L^{\dagger}}(x)$ (or at least the Reeb elevation $\mathcal{D}_{L^{\dagger}}^{\beta}$ of $L^{\dagger}$) in terms of the form $\beta$ in the vicinity of $L^{\dagger}\subset\partial_{1}^{+}X(v_{\beta})$. $\diamondsuit$
@@
Let $\beta$ be a contact form on $X$, and $v_{\beta}$ its traversing boundary generic Reeb vector field. Assume that Property $\mathsf{A}$ from Definition 3.1 is valid. Let $\phi:(L,\partial L)\hookrightarrow(X,\partial_{1}X)$ be a Legendrian embedding of a compact manifold $L$.
@@
  • the Lagrangian immersion $\mathsf{int}(L^{\dagger})\subset\partial_{1}^{+}X(v_{\beta})\setminus(\mathcal{X}_{2}(v_{\beta})\cap\partial_{1}^{+}X(v_{\beta}))$ with respect to the symplectic form $d\beta|_{\mathsf{int}(\partial_{1}^{+}X(v_{\beta}))}$,
  • the set of base points $\{b_{\kappa}\}_{\kappa}$ and the elevation values $\{\psi^{\beta}_{L}(b_{\kappa})\}_{\kappa}$ as in Corollary 7.1,
allow for a reconstruction of the Legendrian embedding $\phi:L\hookrightarrow X$, up to a diffeomorphism of $X$ that is the identity on $\partial_{1}X$.
@@
Let $X$ be a compact connected smooth $2n+1$-dimensional manifold, $\beta$ a contact form, and $v_{\beta}$ its Reeb vector field, which is traversing.
Let $L\subset X$ be a sub-Legendrian $(2k+1)$-dimensional closed submanifold of $X$, where $2k+1\leq n$. We assume that $L$ is in a weakly general position with respect to the waterfall $\mathcal{X}_{2}(v_{\beta})$. We denote by $L(v_{\beta})$ the set of $(-v_{\beta})$-trajectories originating in $L$, and by $\delta L(v_{\beta})$ the set of $(-v_{\beta})$-trajectories originating in $L$ and hitting the locus $\partial_{2}^{+}X(v_{\beta})$ (see Fig. 3).
Then, for any $k>0$, we get $\int_{L(v_{\beta})}(d\beta)^{k+1}=0$ and $\int_{L^{\dagger}}\beta\wedge(d\beta)^{k}=0.$
In contrast, for $k=0$, we still get $\int_{L(v_{\beta})}d\beta=0,\;\text{ but }\;\int_{L^{\dagger}}\beta=-\int_{\delta L(v_{\beta})}\beta\;\leq\;0.$
@@
Note that, at a generic point $x\in L(v_{\beta})$, the tangent space $T_{x}(L(v_{\beta}))$ is spanned by $v_{\beta}(x)$ and the image of a tangent space $T_{y}L$ under $(-v_{\beta})$-flow that takes $y\in L$ to $x$. Then, $(d\beta)^{k+1}|_{T_{x}(L(v_{\beta}))}=0$, since $v_{\beta}\in\ker(d\beta)$ . Therefore, $\int_{L(v_{\beta})}(d\beta)^{k+1}=0$.
@@
In the special case $k=0$, using a similar analysis and Stokes’ theorem, we get
$\int_{L^{\dagger}}\beta=-\int_{\delta L(v_{\beta})}\beta\;\leq\;0$